Engineers are conducting a sea level test of the space shuttle main engine. They are suspicious that the nozzle is not performing up to specifications due to the total pressure losses. To aid in determining its operational characteristics, the engineers have installed a series of wall static pressure taps at the nozzle exit station. Test parameters include: 467 kg/sec of gas leaving the combustion chamber operating at a total pressure and total temperature of 20.7 MPa and 4080 K respectively, and a nozzle exit area equal to 77 times that of the throat area. Assuming the nozzle is running full (without any flow separation) and adiabatically, with a molecular weight of 13.0 and a gamma of 1.25, determine

1. The nozzle throat area (assuming isentropic flow to the throat)

2. The Mach number at the nozzle exit where the average wall static pressure is measured as 0.016MPa

3. The total pressure recovery to the nozzle exit.
Answer #1

I have solved part 1 and 2 for you, i didn't solve the 3 part because the only way to calculate pressure recovery known to me is through the gas table but the gas table of gamma 1.25 is not available to me. if you provide the gas table of gamma 1.25 then i can solve part 3 for you.

throat mlor exit Now, the given parameters are. in = 467 kg/S. P, = 20.7 MPa. T, = 4000K A = 774 å = 0 M =13 8 = 1.25 [ Aez ae [ 2 1.25 +1 1.25 1.25 + 20.7mpa (1.25+1) P = 0.93665 * 20.7 mPa. Pt = 19.3888 MPa с. е. 7 г. - поzz, чJ bя сhооtiq , Now, fputting values in 0 Ą: 467 kg x y 8.314x3626.667 I knop T 30 J 12 molk kg 19.3888x10° Pax $1.25 = 1.03752 K10°m°f the rJJ 4 =T = 973.39 K bon So, velocity of sound at T = 973.39 k ų = 51.4+ 287x973.39 - 625.386 m/s Now, Area at exit Ag = 77*A= 0.0798so, Mach No.- 227.46 625 386 Mach No, M = 0.3637 & Subsonic flow. The stagnation pressure will be [ 2 1.25 5.20.1[14 (1.25-1)section comes to a sudden stop. A normal shock develops at a sechon between the throat & exit plane which tad cause a sudden